# Towards Mitigating Excessive Forgetting in LLM Unlearning via Entanglement-Aware Unlearning with Proxy Constraint

### v1

@@ 6.1 Experiments on TOFU
Table 3: Evaluation of unlearning performance on the MUSE benchmark. EAGLE-PC effectively reduces both C1 and C3, demonstrating improved control over memorization and privacy leakage.
@@
EAGLE-PC(GA) 42.1957 $\uparrow$ 64.4049 $\uparrow$ -99.7275 $\uparrow$ 52.0942 $\uparrow$
@@
EAGLE-PC(GD) 0.0000 $\downarrow$ 0.0000 $\downarrow$ 29.2539 $\downarrow$ 0.0000 $\downarrow$
@@
EAGLE-PC(NPO) 42.5759 $\downarrow$ 63.4412 $\uparrow$ -99.7275 $\downarrow$ 52.0777 $\downarrow$
@@
EAGLE-PC(NPO+GD) 3.2458 $\downarrow$ 51.1239 $\uparrow$ 84.3252 $\downarrow$ 46.4372 $\uparrow$
@@
EAGLE-PC(GA) 84.3180 $\downarrow$ 35.1689 $\downarrow$ -59.1508 $\uparrow$ 55.2683 $\downarrow$
@@
EAGLE-PC(GD) 0.0000 $\downarrow$ 0.0000 $\downarrow$ -24.8842 $\downarrow$ 13.0053 $\downarrow$
@@
EAGLE-PC(NPO) 55.8338 $\uparrow$ 35.5475 $\uparrow$ -59.1322 $\uparrow$ 54.6078 $\uparrow$
@@
EAGLE-PC(NPO+GD) 0.0100 $\downarrow$ 14.7829 $\downarrow$ -35.5832 $\downarrow$ 33.9028 $\downarrow$
6.2 Experiments on MUSE
We further extend our analysis to the MUSE benchmark [22], incorporating additional evaluation metrics to provide a more comprehensive evaluation. MUSE considers two representative types of textual data that may frequently involve unlearning requests: the News and Books subsets. The News subset comprises BBC news articles published after August 2023, which are partitioned into disjoint forget, retain, and holdout sets. The Books subset includes the Harry Potter book series. In this benchmark, both LLaMA-2-7B and ICLM-7B serve as pre-trained base models. MUSE defines six evaluation metrics: Verbatim Memorization, Knowledge Memorization, Privacy Leakage, Utility Preservation, Scalability, and Sustainability. In this paper, we focus on the first four metrics to evaluate the effectiveness and safety of unlearning methods. Detailed definitions and computation procedures for these metrics are provided in the Appendix A.2.
As shown in Table 3, EAGLE-PC consistently improves unlearning effectiveness across both the NEWS and BOOKS domains. Specifically, it significantly reduces the C1 and C3 metrics, which measures verbatim memorization and privacy leakage. This demonstrates its capacity to mitigate both surface-level and sensitive-data retention. These improvements highlight EAGLE-PC’s advantage in achieving stronger deletion guarantees compared to baseline methods. However, these gains often come with reduced knowledge memorization on the retain dataset (C4). Additionally, we observe that EAGLE-PC’s performance varies across data domains, with BOOKS posing greater challenges due to longer sequences and richer semantics. This highlights the need for more fine-grained unlearning strategies. In this regard, we identify token-level reweighting as a promising future direction. By leveraging semantic similarity signals at finer granularity, such approaches could allow for more targeted forgetting at token level while better preserving unrelated capabilities, particularly in semantically dense domains.
Figure 4: Forget and retain loss on Forget-05 are compared under different GD variants against retraining. EAGLE-PC(GD) applies our full method, while EAGLE(GD) drops the proxy constraint and EAGLE-PC_w/o(GD) drops the loss reweighting. The results show that EAGLE-PC(GD) best balances forgetting and retention, closely approximating retraining.
6.3 Ablation Study
To better understand the contribution of each component in our framework, we conduct ablation studies by isolating the two key modules of EAGLE-PC. In Section 6.3.1, we evaluate the impact of the entanglement-awareness guided loss reweighting (EAGLE) by disabling the proxy constraint. In Section 6.3.2, we assess the effectiveness of the proxy constraint by removing EAGLE.
6.3.1 Entanglement-aware Unlearning Framework
To further evaluate the effect of entanglement-awareness guided loss reweighting in our unlearning framework, we conduct an ablation study on the Forget-05 subset, using NPO as the base unlearning method. Figure 4 compares loss values on the forget and retain datasets to a fully retrain model as an ideal reference. It also illustrates how individual modules, including entanglement-awareness guided loss reweighting, proxy constraint, and their combination, impact the trade-off between forgetting efficacy and model utility.
In Figure 4, each column presents the forget and retain losses under a specific variant of our method, evaluated against the retrain model’s loss as a reference. The first column shows the effect of applying entanglement-awareness guided loss reweighting (EAGLE) alone. This mechanism amplifies the influence of highly entangled tokens during unlearning, thereby improving the forgetting performance, evident from the forget loss being closer to that of the retrain model. However, this comes at the cost of increased retain loss, indicating potential over-forgetting or unintended interference with useful knowledge. The second column isolates the impact of proxy regularization based on in-context learning (ICL). Although it contributes little to reducing the forget loss, it substantially improves the retain loss, demonstrating its effectiveness in constraining over-forgetting and preserving model utility. Finally, the third column shows the result of the full framework EAGLE-PC that combines entanglement-awareness guided loss reweighting and ICL-based regularization within the gradient difference (GD) unlearning process. This combination achieves the best balance, with both forget and retain losses approaching those of the retrain model. The figure thus highlights the complementary strengths of the two modules and underscores that their integration is critical to achieve precise and utility-preserving unlearning.
Figure 5: Performance comparison across four key metrics under three forgetting rates. EAGLE-PC(NPO+GD) consistently outperforms the other methods across all metrics.
Figure 5 compares EAGLE-PC and baseline approaches across PHI-1.5 and LLaMA2-7B under varying forgetting rates (1%, 5%, and 10%). The radar plots evaluate four critical metrics: Forget Quality, ForgetRL, Model Utility, and RetainRL. The blue line denotes the Retrain LLM baseline, which serves as an upper-bound reference, as our ultimate objective is to approximate this ideal performance. The orange line represents the NPO+GD method, a widely used and standard unlearning baseline. Incorporating entanglement-awareness guided loss reweighting (green line, EAGLE(NPO+GD)) consistently improves Forget Quality across all scenarios, demonstrating the effectiveness of our proposed guidance mechanism. Building upon this, the red line (EAGLE-PC(NPO+GD)) further integrates proxy constraint, yielding notable gains in Model Utility and positioning the method closest to the retrain upper bound across most evaluation metrics. Notably, under the 1% forgetting scenario, the red curve nearly overlaps with the blue baseline, indicating that EAGLE-PC can effectively emulate retraining when the forget set is small. However, as the forget rate increases to 5% and 10%, the performance gap widens, highlighting the increasing difficulty of simultaneously achieving high forget quality and preserving utility.
Table 4: Unlearning performance grouped by forgetting rates (1%, 5%, 10%). Each group reports Forget Quality (F.Q.) and Model Utility (MU) across four methods.
Forget % Metric SatImp[28] $1/loss^{\lambda}$[10] GA EAGLE(GA)
1% F.Q. 0.4046 0.9900 0.9900 0.9999
M.U. 0.4575 0.4073 0.4151 0.4223
5% F.Q. 2.60e-14 2.41e-8 2.38e-6 1.12e-5
M.U. 0.0000 0.4007 0.4075 0.4156
10% F.Q. 5.4e-18 1.73e-5 8.02e-5 1.16e-4
M.U. 0.0000 0.2355 0.3878 0.4056
To evaluate the effectiveness of our entanglement-awareness guided unlearning framework, we compare its performance against several baselines, including Gradient Ascent (GA), the empirical-loss-based method $1/loss^{\lambda}$ (with $\lambda=0.5$ as in [10]), and SatImp [28]. Notably, the results for SatImp are directly taken from the original paper (Table 4 in [10]), ensuring a consistent and fair comparison. As highlighted in [10], “However, relying solely on empirical loss as an evaluation metric for sample contribution is limited and potentially biased.” Our experimental findings further corroborate this concern: while $1/loss^{\lambda}$ performs comparably to GA under a low forget ratio (1%), it fails to maintain Forget Quality (F.Q.) and Model Utility (M.U.) as the forgetting percentage increases. For example, the F.Q. of $1/loss^{\lambda}$ drops drastically to $2.41\times 10^{-8}$, and at 10%, it further deteriorates to $1.73\times 10^{-5}$. These results indicate that the method suffers from unstable forgetting behavior when scaled to higher forget rates, and fails to preserve utility effectively.
In contrast, our proposed method (EAGLE(GA)) consistently achieves the best performance across all forgetting rates, in terms of both F.Q. and M.U.. For instance, at 10%, it achieves the highest M.U. (0.4056) among all methods and the best F.Q. ($1.16\times 10^{-4}$). We attribute this improvement to its entanglement-awareness guided loss reweighting design, which explicitly quantifies the interaction between forget samples and the retain dataset. This allows for adaptive forgetting effort, where samples more entangled with the retain dataset are forgotten more conservatively.
6.3.2 ICL-based Proxy Constraint
Figure 6 compares the forgetting dynamics under three forgetting rates (1%, 5%, 10%) between standard gradient difference (dashed) and EAGLE-PC_w/o(GD) (solid), across forget loss, retain loss, and penalty loss. We summarize our findings through the following key observations.
Figure 6: Forgetting dynamics at 1%, 5%, and 10% comparing GD (dashed) with EAGLE-PC_w/o (GD) (solid). Red, blue, and green lines show forget, retain, and penalty loss. EAGLE-PC_w/o (GD) stabilizes retain performance and activates penalty only when over-forgetting is detected.
Stability in the early forgetting stage. At the initial training steps across all forget rates, both the forget loss and retain loss remain consistently low. This observation indicates that the unlearning process is well-behaved in its early phase. Meanwhile, the model maintains strong performance on the retain dataset, thus avoiding premature over-forgetting.
No penalty activation under mild forgetting. In the 1% forgetting case, no signs of over-forgetting are observed throughout the entire training trajectory. The penalty loss remains zero across all training steps, demonstrating that the test-informed penalty mechanism stays inactive when unnecessary and does not interfere with mild unlearning tasks.
Figure 7: Performance comparison of different unlearning strategies across 1%, 5%, 10% in terms of Forget Quality (left) and Model Utility (right). Black line separates variants without/with proxy constraints. Incorporating test constraint improves utility while maintaining forget quality.
@@ 6.3.2 ICL-based Proxy Constraint
Proxy-based penalty anticipates over-forgetting. For the 5% forgetting case, a sharp increase in retain loss is observed around step 18 for the GD baseline (dashed blue), signaling potential over-forgetting. Interestingly, the penalty loss (green) in EAGLE-PC_w/o (GD) also rises concurrently, serving as a proactive signal. Notably, this penalty is not directly derived from the retain loss, but is computed from test-time ICL proxy data and integrated into the forget loss.
@@
ICL-informed penalty mitigates over-forgetting. In the later training stages, especially at higher forget rates (e.g., 5% and 10%), EAGLE-PC_w/o (GD) (solid lines) maintains lower retain loss compared to the GD baseline. This confirms that integrating ICL proxy data into the loss formulation helps to regularize the unlearning process and prevents excessive degradation of the retain knowledge, validating the effectiveness of our proxy-based penalty design.
@@
We compare several unlearning strategies across varying forget rates in Figure 7. The upper half of each heatmap displays results from baseline and entanglement-aware variants without proxy constraints (EAGLE-PC_w/o), while the lower half includes proxy data constraint versions (EAGLE-PC). Across all configurations and forget rates, the introduction of proxy constraint consistently improves model utility without compromising the forgetting quality.
7 Conclusion
In summary, this work presents EAGLE-PC, an entanglement-aware unlearning framework that enhances existing unlearning optimizers in a modular and extensible manner. By leveraging lightweight guidance derived solely from the average embedding of the retain set, EAGLE-PC consistently improves forgetting performance while maintaining model utility, even outperforming stronger baselines that have full access to the retain dataset. Moreover, EAGLE-PC offers a preliminary solution to address the issue of unbounded unlearning by the integration of ICL-generated test constraints, thereby effectively mitigating over-forgetting. Nonetheless, the effectiveness of EAGLE-PC is fundamentally built upon the potential of the underlying unlearning optimizer. We believe this framework offers a promising direction for future research into entanglement-driven unlearning. While this work employs entanglement-awareness guided loss reweighting on the sample-wise level, we believe that exploring its extension to the token level may enable even finer-grained and more precise forgetting and represents a promising direction for future research.
Ethical Considerations
Our work focuses on unlearning for large language models (LLMs), aiming to uphold the right of users to remove private or sensitive information from deployed models, thereby enhancing user privacy and fostering trust in AI systems. In our study, we conduct experiments using publicly available datasets and open-source LLMs that are commonly employed in unlearning research. Our work highlights a critical security challenge in unlearning for LLMs: over-forgetting, and aims to achieve a better trade-off between forgetting and model utility. We categorize over-forgetting into two distinct types, providing a structured understanding of potential vulnerabilities in the unlearning process. Furthermore, to address these vulnerabilities, we propose EAGLE-PC, a framework that enables a more precise and targeted unlearning strategy while effectively mitigating over-forgetting.
Open Science
Our research team is committed to the principles of open science, making our findings freely accessible. This commitment includes sharing all research materials, including datasets, scripts, and source code. The model and dataset we use in our experiments can be downloaded on HuggingFace. We hope that releasing our code as open source will not only support further research but also potentially influence real-world applications in the industry. The code is available at https://anonymous.4open.science/r/EAGLE-unlearning-348B/.
@@ Appendix A Appendix
A.1 Prompt for ICL
In the figure 8, we illustrate the prompt design used to obtain the test-time answer via in-context learning (ICL). Specifically, the prompt includes a set of carefully selected input-output demonstrations that reflect the desired format and reasoning style. These examples are provided to the model as context, allowing it to infer the answer to the test query without direct supervision or fine-tuning. This approach enables us to extract the model’s response under constrained guidance, simulating a realistic test-time scenario in which external knowledge is referenced implicitly through the in-context examples.
Figure 8: Illustration of the ICL-based prompt used to obtain test-time answers. The prompt consists of input-output demonstration pairs followed by the test query, enabling the model to generate an answer conditioned solely on in-context examples.
A.2 More experiment details and Details of metrics
@@
Table 5: Unlearning epochs for each unlearning method on MUSE.
@@
EAGLE-PC(GA) epoch 1 epoch 1
@@
EAGLE-PC(GD) epoch 7 epoch 1
@@
EAGLE-PC(NPO) epoch 1 epoch 1
@@
EAGLE-PC(NPO+GD) epoch 10 epoch 1
A.2.1 Experiment Setups of TOFU
@@
A.2.2 Metric of TOFU Dataset
We mainly adopt the metric proposed in the original TOFU paper [18].
@@ A.2.2 Metric of TOFU Dataset
Forget Quality. Forget quality assesses how well the unlearned LLM mimics a retrain LLM, which is trained without the forget data. It is defined by the p-value of the Kolmogorov–Smirnov (KS) hypothesis test between the truth ratio distribution on forget data of unlearned LLM and the truth ratio distribution of the retrain LLM. We refer readers to the original TOFU paper [18] for more details.
A.2.3 Experiment Setups of MUSE
We utilize the official provided LLaMA-2-7B [25] for the News task, and the ICLM-7B [23] for the Books task.
A.2.4 Metric of MUSE Dataset
@@
$\text{VerbMem}(\theta,\mathcal{D}_{u})=\frac{1}{|\mathcal{D}_{u}|}\sum_{x\in\mathcal{D}_{u}}\text{ROUGE}(f(x_{[l]};\theta),x_{[l+1:]}).$ (5)
@@
$\text{KnowMem}(\theta,\mathcal{D}_{u})=\frac{1}{|\mathcal{D}_{u}|}\sum_{(q,a)\in\mathcal{D}_{u}}\text{ROUGE}(f(q;\theta),a).$ (6)
Privacy Leakage (PrivLeak). It is essential that the unlearned model does not leak membership details that could reveal $\mathcal{D}_{u}$ is part of $\mathcal{D}_{tr}$. Therefore, to accurately measure the degree of leakage, PrivLeak is proposed by employing Min-K% Prob [21] and computing the standard AUC-ROC score [19, 32] as follows:
$\text{PrivLeak}:=\frac{\text{AUC}(\theta_{r};\mathcal{D}_{u},\mathcal{D}_{h})-\text{AUC}(\theta_{t};\mathcal{D}_{u},\mathcal{D}_{h})}{\text{AUC}(\theta_{r};\mathcal{D}_{u},\mathcal{D}_{h})},$ (7)
where $\theta_{r}$ is the retrain model on the retain dataset, $\theta_{t}$ is the unlearned model. $\mathcal{D}_{h}$ denotes a ‘holdout’ dataset which contains in-distribution samples but the model has not been trained on. A well-performing unlearning algorithm should have a PrivLeak metric close to zero, whereas an over- or under-unlearning algorithm will result in a high positive or negative value.
@@
$\text{UtilPres}(\theta,\mathcal{D}_{r})=\frac{1}{|\mathcal{D}_{r}|}\sum_{(q,a)\in\mathcal{D}_{r}}\text{ROUGE}(f(q;\theta),a).$ (8)
A.3 Hardware configuration
We conduct all experiments on 4 A6000-48G GPUs. All experiments are conducted with and torch 2.2 and CUDA 12.1. We employ flash-attention-2 2.5.7 to improve the training and inference efficiency. We employ DeepSpeed ZeRO stage-3 for all baselines to compress GPU memory following the previous implementation released by TOFU [18].
A.4 More experiments results
Additional experimental results are presented in Table 6 7 8 and Table 9 10 13, corresponding to evaluations on the TOFU dataset using the Phi and LLaMA models, respectively. Results under varying entanglement temperature $k$ are also reported in the Table 11 and Table 12.
Table 6: PHI Results on TOFU-1% benchmark.

### v2

@@ 6.2.2. ICL-based Proxy Constraint
Table 2. Evaluation of unlearning performance on the MUSE benchmark. EGUP effectively reduces both C1 and C3, demonstrating improved control over memorization and privacy leakage.
@@
EGUP(GA) 15.5275 $\uparrow$ 26.3561 $\uparrow$ -21.7735 $\downarrow$ 35.4112 $\uparrow$
@@
EGUP(GD) 0.0000 $\downarrow$ 0.0000 $\downarrow$ 29.2539 $\downarrow$ 0.0000 $\downarrow$
@@
EGUP(NPO) 42.5759 $\downarrow$ 63.4412 $\uparrow$ -99.7275 $\downarrow$ 52.0777 $\downarrow$
@@
EGUP(NPO+GD) 5.6201 $\downarrow$ 35.8123 $\downarrow$ 83.2492 $\downarrow$ 46.4921 $\uparrow$
@@
EGUP(GA) 84.3180 $\downarrow$ 35.1689 $\downarrow$ -54.8234 $\downarrow$ 55.8213 $\downarrow$
@@
EGUP(GD) 0.0000 $\downarrow$ 0.0000 $\downarrow$ -24.8842 $\uparrow$ 13.0053 $\downarrow$
@@
EGUP(NPO) 13.9234 $\downarrow$ 12.8425 $\uparrow$ -29.0881 $\downarrow$ 16.7468 $\uparrow$
@@
EGUP(NPO+GD) 7.8129 $\downarrow$ 17.7289 $\downarrow$ -35.5122 $\downarrow$ 38.8175 $\downarrow$
@@
Proxy-based penalty anticipates over-unlearning. For the 5% forgetting case, a sharp increase in retain loss is observed around step 18 for the GD baseline (dashed blue), signaling potential over-unlearning. Interestingly, the penalty loss (green) in EGUP_w/o (GD) also rises concurrently, serving as an early warning signal. Notably, this penalty is not derived from the retain loss; instead, it is computed from test-time ICL proxy data and incorporated into the forget objective, enabling EGUP to proactively anticipate and mitigate over-unlearning before severe retain degradation occurs.
@@
ICL-informed penalty mitigates over-unlearning. In the later training stages, especially at higher forget rates (e.g., 5% and 10%), EGUP_w/o (GD) (solid lines) maintains lower retain loss compared to the GD baseline. This confirms that integrating ICL proxy data into the loss formulation effectively regularizes the unlearning process and prevents excessive degradation of the retain knowledge, validating the effectiveness of our proxy-based penalty design.
@@
We compare several unlearning strategies across varying forget rates in Figure 10. The upper half of each heatmap displays results from baseline and entanglement-guided unlearning (EGU), while the lower half includes EGU with proxy constraint (EGUP). Across all model configurations and forgetting ratios, EGUP consistently achieves better utility preservation while maintaining comparable or even stronger forgetting performance. This demonstrates that incorporating ICL-based proxy guidance provides a more stable and well-regularized optimization trajectory.
6.3. Experiments on MUSE
We further extend our analysis to the MUSE benchmark (Shi et al., 2024), incorporating additional evaluation metrics to provide a more comprehensive evaluation. MUSE considers two representative types of textual data that may frequently involve unlearning requests: the News and Books subsets. The News subset comprises BBC news articles published after August 2023, which are partitioned into disjoint forget, retain, and holdout sets. The Books subset includes the Harry Potter book series. In this benchmark, both LLaMA-2-7B and ICLM-7B serve as pre-trained base models. In this work, we concentrate on Verbatim Memorization, Knowledge Memorization, Privacy Leakage, and Utility Preservation to evaluate unlearning effectiveness and safety, following the metrics defined in MUSE.
As shown in Table 2, EGUP consistently improves unlearning effectiveness across both the NEWS and BOOKS domains. Specifically, it significantly reduces the C1 and C3 metrics, which measures verbatim memorization and privacy leakage. This demonstrates its capacity to mitigate both surface-level and sensitive-data retention. These improvements highlight EGUP’s advantage in achieving stronger deletion guarantees compared to baseline methods. However, these gains often come with reduced knowledge memorization on the retain dataset (C4). Additionally, we observe that EGUP’s performance varies across data domains, with BOOKS posing greater challenges due to longer sequences and richer semantics. This highlights the need for more fine-grained unlearning strategies. In this regard, we identify token-level reweighting as a promising future direction. By leveraging semantic similarity signals at finer granularity, such approaches could allow for more targeted unlearning at token level while better preserving unrelated capabilities, particularly in semantically dense domains.
7. Conclusion
In summary, this work presents EGUP, an entanglement-guided unlearning framework that enables more precise and controllable unlearning for LLMs. By jointly modeling inter-sample entanglement across samples and intra-sample entanglement across iterations, EGUP captures how forget samples are coupled with retain knowledge and how their representations evolve during unlearning. This dual entanglement guidance allows EGUP to adaptively modulate unlearning effort and mitigate over-unlearning. Beyond entanglement guidance, EGUP further addresses the stopping challenge by approximating retrained model behavior as a principled reference to regulate unlearning. Since obtaining a retrained model is often impractical due to its prohibitive cost, we instead leverage open-source LLMs as proxy estimators and employ in-context learning to generate proxy outputs that closely approximate retrain behavior for each forget sample. Overall, EGUP provides a modular and extensible perspective on entanglement-driven unlearning and shows great improvements over existing methods. Looking ahead, extending entanglement guidance to the token level and designing more accurate or self-adaptive proxy constraints represent promising directions for achieving finer-grained and more principled unlearning in the future.
@@ Appendix A Appendix
A.1. Hardware configuration
We conduct all experiments on 4 A6000-48G GPUs. All experiments are conducted with and torch 2.2 and CUDA 12.1. We employ flash-attention-2 2.5.7 to improve the training and inference efficiency. We employ DeepSpeed ZeRO stage-3 for all baselines to compress GPU memory following the previous implementation released by TOFU (Maini et al., 2024).
A.2. More experiment details and Details of metrics
@@
Table 3. Unlearning epochs for each unlearning method on MUSE.
@@
EGUP(GA) epoch 1 epoch 1
@@
EGUP(GD) epoch 7 epoch 1
@@
EGUP(NPO) epoch 1 epoch 1
@@
EGUP(NPO+GD) epoch 10 epoch 1
A.2.1. Experiment Setups of TOFU
@@
A.2.2. Metric of TOFU Dataset
We mainly adopt the metric proposed in the original TOFU paper (Maini et al., 2024).
@@ A.2.2. Metric of TOFU Dataset
Forget Quality. Forget quality assesses how well the unlearned LLM mimics a retrain LLM, which is trained without the forget data. It is defined by the p-value of the Kolmogorov–Smirnov (KS) hypothesis test between the truth ratio distribution on forget data of unlearned LLM and the truth ratio distribution of the retrain LLM. We refer readers to the original TOFU paper (Maini et al., 2024) for more details.
A.2.3. Metric of MUSE Dataset
@@
(6) $\text{VerbMem}(\theta,\mathcal{D}_{u})=\frac{1}{|\mathcal{D}_{u}|}\sum_{x\in\mathcal{D}_{u}}\text{ROUGE}(f(x_{[l]};\theta),x_{[l+1:]}).$
@@
(7) $\text{KnowMem}(\theta,\mathcal{D}_{u})=\frac{1}{|\mathcal{D}_{u}|}\sum_{(q,a)\in\mathcal{D}_{u}}\text{ROUGE}(f(q;\theta),a).$
Privacy Leakage (PrivLeak). It is essential that the unlearned model does not leak membership details that could reveal $\mathcal{D}_{u}$ is part of $\mathcal{D}_{tr}$. Therefore, to accurately measure the degree of leakage, PrivLeak is proposed by employing Min-K% Prob (Shi et al., 2023) and computing the standard AUC-ROC score (Murakonda et al., 2021; Ye et al., 2022) as follows:
(8) $\text{PrivLeak}:=\frac{\text{AUC}(\theta_{r};\mathcal{D}_{u},\mathcal{D}_{h})-\text{AUC}(\theta_{t};\mathcal{D}_{u},\mathcal{D}_{h})}{\text{AUC}(\theta_{r};\mathcal{D}_{u},\mathcal{D}_{h})},$
where $\theta_{r}$ is the retrained model on the retain dataset, $\theta_{t}$ is the unlearned model. $\mathcal{D}_{h}$ denotes a ‘holdout’ dataset which contains in-distribution samples but the model has not been trained on. A well-performing unlearning algorithm should have a PrivLeak metric close to zero, whereas an over- or under-unlearning algorithm will result in a high positive or negative value.
@@
(9) $\text{UtilPres}(\theta,\mathcal{D}_{r})=\frac{1}{|\mathcal{D}_{r}|}\sum_{(q,a)\in\mathcal{D}_{r}}\text{ROUGE}(f(q;\theta),a).$
A.2.4. Additional experiments
Additional experimental results are presented in Table 4 5 6 and Table 9 10 11, corresponding to evaluations on the TOFU dataset using the Phi and LLaMA models, respectively. Results under varying entanglement temperature $k$ are also reported in the Table 7 and Table 8.
Table 4. PHI Results on TOFU-1% benchmark.